OpenStudy (anonymous):

if you divide 9x^3+3x^2-6x-3 to 3x-1 using synthetic division, how it is going to solve?

OpenStudy (anonymous):

\[\frac{9x^3+3x^2-6x-3}{3x-1}\] You can't use synthetic division right away, because the divisor (the bottom part) has to be in the format of \(x-k\), without a number out in front of it. First factor out a 3 from the denomenator: \[\frac{9x^3+3x^2-6x-3}{3(x-1)} = \frac{1}{3}\times\frac{9x^3+3x^2-6x-3}{x-\dfrac{1}{3}}\] Now try synthetic division.

OpenStudy (anonymous):

9 6 \[\frac{ -4 }{ 3 }\] i got this answer and i have no idea what to do next because i encouter a fraction. HOw is it going to be?

OpenStudy (anonymous):

Well you can first cancel the \(\large^1/_3\) and the numerator to \(\large3x^3+x^2-2x-1\). Then divide:\[ \large{\left.\underline{^1/_3}\right\vert~~~~~~3~~~~1~~ -2 ~ ~~-1 \\ ~~~~~~~~~~~~~~~~~~~1~~^{-2}/_3~~~^{-8}/_9 \\ ~~~~~~~~~~~~~\overline{3~~~~2~~^{-8}/_3~~\left\vert^{-11}/_3\right.}}\] So\[3x^2 + 2x - \frac{8}{3} - \frac{\dfrac{11}{3} }{x - \dfrac{1}{3} }\]

OpenStudy (kropot72):

 so \[3x ^{2}+2x-\frac{4}{3}-\frac{\frac{13}{9}}{x-\frac{1}{3}}\]

OpenStudy (anonymous):

Oh yeah, true 2/3 = \(\dfrac{2}{3}\), not \(-\dfrac{2}{3}\)